OpenStudy (countrygirl1431):

Write the standard form of the line that has a slope of - 3/4and y-intercept of -2. Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (countrygirl1431):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you know slope intercept form?

OpenStudy (anonymous):

\[y=mx+b\]

OpenStudy (countrygirl1431):

yeah i know that one by heart

OpenStudy (countrygirl1431):

@surjithayer

OpenStudy (anonymous):

m=slope, b=y intercept

OpenStudy (countrygirl1431):

so would it be y=-3/4x+-2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how can we get rid of the fraction?

OpenStudy (countrygirl1431):

multiply by 4

jimthompson5910 (jim_thompson5910):

multiply both sides by 4, yep

jimthompson5910 (jim_thompson5910):

so y=-3/4x-2 will turn into 4y = -3x-8 after you multiply both sides by 4

jimthompson5910 (jim_thompson5910):

what comes next?

OpenStudy (countrygirl1431):

add 3x to both sides

jimthompson5910 (jim_thompson5910):

yep to get 3x+4y = -8

OpenStudy (countrygirl1431):

thanks

jimthompson5910 (jim_thompson5910):

np